# Automated Federated Pipeline for Parameter-Efficient Fine-Tuning of Large Language Models

### v1

@@ 5.1. Overall Performance Evaluation
Table 3. GPT-2 large model with different combinations of trainable weights on the E2E NLG challenge.
@@
Table 4. GPT-2 large model with different heterogeneous configurations on the E2E NLG challenge.
@@
Table 5. FedPipe on GPT-2 large model with different quantization bits on the E2E NLG challenge.
5.2. Ablation Evaluation
Important weights identification (IWI). Figure LABEL:fig:hetero_weights shows the impact of the selection of trainable weights on GPT2-L model under heterogeneous setting. Considering that the average number of trainable weights in our training process is approximately 3, we compare the IWI method with a random combination of three trainable weights for fair comparison. The proposed IWI method exhibits lower PPL than all three random combinations of weights, which demonstrates the effectiveness of our IWI method. This is because large singular values usually correspond to the primary directions of variation in model weights, reflecting where the density of principal information is high. Therefore, it is efficient to prioritize fine-tuning the trainable weights with large singular values. Additionally, training performance on other metrics is summarized in Table 3.
Heterogeneous configuration (HC). Figure LABEL:fig:hetero_avg_var_gpt2 presents the impact of the selection of rank size on model training. We compare our proposed HC method with conventional LoRA schemes with different fixed rank sizes. It is observed that as the rank increases from 4 to 16, the PPL of the conventional LoRA framework grows rapidly from 3.23 to 3.51, nearly 9%, indicating an increasingly severe influence of the heterogeneous computing resource across clients on model training. This is attributed to the observation that the larger rank implies heavier computing workloads, leading to lower under-training rates. Conversely, the PPL of the HC method increases slightly with the rank, owing to our design of adaptively adjusting the rank based on heterogeneous computing capabilities. Moreover, the performance gap between HC method and LoRA with fixed rank sizes significantly widens with the increase in rank, further revealing the effectiveness of HC method. Similarly, the training performance on other metrics is shown in Table 4.
Quantization. Figure LABEL:fig:GPT2_lg_quantization illustrates the GPU memory space and training accuracy under different quantization bits. The GPU memory space of FP32 is significantly larger than other quantization levels, 20.5 GB, where the model takes up 3.14 GB, which is approximately 1.9, 3.2, and 4.9 times that of FP16, NF8, and NF4, respectively. Quantization bit makes trade off between memory space and training accuracy of LLMs. It is observed that as the number of quantization bits diminishes, the training accuracy of the model only slightly decreases. FP32 achieves slightly higher training accuracy, i.e., with PPL 3.224, whereas FP16, NF8, and NF4 exhibit similar lower training accuracy, with PPL approximately 3.228. The proposed quantization scheme further reduces the memory requirements of the FedPipe framework, thereby improving its scalability. A more comprehensive comparison is illustrated in Table 5.
6. Related Work
@@
Parameter-efficient fine-tuning: PEFT has recently emerged as a promising approach for efficient LLM fine-tuning, which strikes a balance between the number of trainable parameters and training accuracy in NLP tasks. Among various PEFT approaches, ADAPTER (Houlsby et al., 2019; Pfeiffer et al., 2021; Karimi Mahabadi et al., 2021) and LoRA (Hu et al., 2022) stand out as the most attractive techniques for substantially reducing trainable parameters without compromising LLM fine-tuning performance. ADAPTER inserts small trainable modules into each transformer block while keeping the parameters of the pre-trained model frozen to improve training efficiency. Compacter (Karimi Mahabadi et al., 2021) is a variant of ADAPTER that achieves lower parameter complexity based on shared information across adapters and low-rank subspaces of the model. However, the ADAPTER approaches introduce additional computing overhead in adapter layers, increasing the LLM fine-tuning latency, especially with smaller batch sizes. In contrast, LoRA (Hu et al., 2022) optimizes the variation of dense layers with low-rank decomposition matrices while keeping pre-trained model parameters frozen. Unlike ADAPTER, the low-rank matrices are inserted into the fine-tuned modules in a parallel manner, mitigating any additional training latency and thus enhancing the efficiency of LLM fine-tuning.
@@
FL paradigm for LLMs: PEFT needs to fine-tune the LLMs based on domain-specific data of downstream tasks. However, clients may be reluctant to share raw data with the server due to data privacy concerns, especially privacy-sensitive data such as medical images. A promising solution to this issue is the FL paradigm, which enables multiple users to train models without sharing raw data. Recently, some research efforts have been made to merge LLM fine-tuning with the FL paradigm.  (Lin et al., 2022) is the first work to integrate FL with LLM fine-tuning, comprehensively validating the effectiveness of the FedNLP framework on four common formulations of NLP tasks: text classification, sequence tagging, question answering, and seq2seq generation. FedAdapter (Cai et al., 2023b) addresses the high training cost of FedNLP, which expedites the model convergence rate by progressively upgrading adapter configurations as well as continuously profiling future adapter configurations. However, existing works only focus on the traditional language models such as BART or BERT, whereas the most prevalent LLMs (e.g., GPT-2 (Radford et al., 2019)) with a significant number of parameters are ignored.
@@
Model compression and quantization: The unprecedented number of parameters for LLMs renders computation and storage significant bottlenecks for LLM fine-tuning. Model compression and quantization have emerged as promising techniques to reduce model storage size and accelerate inference by decreasing bit precision (Bernstein et al., 2018; Wu et al., 2018). GPTQ (Frantar et al., 2023) leverages second-order information for error compensation, significantly enhancing model accuracy and training efficiency. The authors of (Lin et al., 2023c) propose an edge-cutting quantization method, named AWQ, to quantize weights into low-bit integers, thereby reducing the hardware barrier and speeding up token generation. Unlike previous work on quantization for LLM inference, SwitchBack layers (Wortsman et al., 2023) investigates backpropagation through quantized weights over 1B parameters and QLoRA (Dettmers et al., 2023) quantizes weights into 4-bit NormalFloat and backpropagates gradients into low-rank structure with the pre-trained LLMs.
@@
Resource heterogeneity: The resource heterogeneity (Horvath et al., 2021; Lin et al., 2023a; Li et al., 2022) poses a significant challenge to deploying FL frameworks (Wang et al., 2021; Tran et al., 2019; Wang et al., 2018) for LLMs. Clients with limited computing or memory resources require more time to complete local model updates, impeding the FL training process or even being excluded from model training. To the best of our knowledge, there is no existing work that addresses resource heterogeneity in FL frameworks for LLM fine-tuning.
7. Conclusion
@@
In this paper, we have proposed, and implemented FedPipe, an automated federated pipeline, to facilitate LLM FL fine-tuning in edge servers with heterogeneous computing resources. We have modeled the problem as a MILP to guide our design. To solve the problem, we have first selected important weights, and configured the corresponding parameters under computing budgets of edge servers in their LoRA adapters. We have also quantized LoRA adapters to further reduce memory space. Our extensive evaluations have demonstrated that FedPipe achieves better performance than the state-of-the-art baselines in LLM FL.
This work has demonstrated the potential of integrating FL and LLM PEFT paradigm. However, the staggering model size of LLMs imposes heavy computing and communication burdens on clients, posing significant barriers to the democratization of the FL LLM fine-tuning paradigm. Therefore, the integration of split learning (Lin et al., 2024a; Vepakomma et al., 2018; Lin et al., 2024b, c) and LLM PEFT paradigms is worth further exploration.

### v2

@@ V-B Ablation Evaluation
TABLE III: GPT-2 large model with different combinations of trainable weights on the E2E NLG challenge.
Heterogeneous configuration (HC). Fig. 16b presents the impact of the selection of rank size on model training. We compare our proposed HC method with conventional LoRA schemes with different fixed rank sizes. It is observed that as the rank increases from 4 to 16, the PPL of the conventional LoRA framework grows rapidly from 3.23 to 3.51, nearly 9%, indicating an increasingly severe influence of the heterogeneous computing resource across clients on model training. This is attributed to the observation that the larger rank implies heavier computing workloads, leading to lower under-training rates. Conversely, the PPL of the HC method increases slightly with the rank, owing to our design of adaptively adjusting the rank based on heterogeneous computing capabilities. Moreover, the performance gap between HC method and LoRA with fixed rank sizes significantly widens with the increase in rank, further revealing the effectiveness of HC method. Similarly, the training performance on other metrics is shown in Table IV.
@@
TABLE IV: GPT-2 large model with different heterogeneous configurations on the E2E NLG challenge.
Quantization. Fig. 16c illustrates the GPU memory space and training accuracy under different quantization bits. The GPU memory space of FP32 is significantly larger than other quantization levels, 20.5 GB, where the model takes up 3.14 GB, which is approximately 1.9, 3.2, and 4.9 times that of FP16, NF8, and NF4, respectively. Quantization bit makes trade off between memory space and training accuracy of LLMs. It is observed that as the number of quantization bits diminishes, the training accuracy of the model only slightly decreases. FP32 achieves slightly higher training accuracy, i.e., with PPL 3.224, whereas FP16, NF8, and NF4 exhibit similar lower training accuracy, with PPL approximately 3.228. The proposed quantization scheme further reduces the memory requirements of the FedPipe framework, thereby improving its scalability. A more comprehensive comparison is illustrated in Table V.
@@
TABLE V: FedPipe on GPT-2 large model with different quantization bits on the E2E NLG challenge.
VI Related Work
@@
Parameter-efficient fine-tuning: PEFT has recently emerged as a promising approach for efficient LLM fine-tuning, which strikes a balance between the number of trainable parameters and training accuracy in NLP tasks. Among various PEFT approaches, ADAPTER [35, 36, 38] and LoRA [37] stand out as the most attractive techniques for substantially reducing trainable parameters without compromising LLM fine-tuning performance. ADAPTER inserts small trainable modules into each transformer block while keeping the parameters of the pre-trained model frozen to improve training efficiency. Compacter [38] is a variant of ADAPTER that achieves lower parameter complexity based on shared information across adapters and low-rank subspaces of the model. However, the ADAPTER approaches introduce additional computing overhead in adapter layers, increasing the LLM fine-tuning latency, especially with smaller batch sizes. In contrast, LoRA [37] optimizes the variation of dense layers with low-rank decomposition matrices while keeping pre-trained model parameters frozen. Unlike ADAPTER, the low-rank matrices are inserted into the fine-tuned modules in a parallel manner, mitigating any additional training latency and thus enhancing the efficiency of LLM fine-tuning.
@@
FL paradigm for LLMs: PEFT needs to fine-tune the LLMs based on domain-specific data of downstream tasks. However, clients may be reluctant to share raw data with the server due to data privacy concerns, especially privacy-sensitive data such as medical images. A promising solution to this issue is the FL paradigm, which enables multiple users to train models without sharing raw data. Recently, some research efforts have been made to merge LLM fine-tuning with the FL paradigm.  [65] is the first work to integrate FL with LLM fine-tuning, comprehensively validating the effectiveness of the FedNLP framework on four common formulations of NLP tasks: text classification, sequence tagging, question answering, and seq2seq generation. FedAdapter [43] addresses the high training cost of FedNLP, which expedites the model convergence rate by progressively upgrading adapter configurations as well as continuously profiling future adapter configurations. However, existing works only focus on the traditional language models such as BART or BERT, whereas the most prevalent LLMs (e.g., GPT-2 [2]) with a significant number of parameters are ignored.
@@
Model compression and quantization: The unprecedented number of parameters for LLMs renders computation and storage significant bottlenecks for LLM fine-tuning. Model compression and quantization have emerged as promising techniques to reduce model storage size and accelerate inference by decreasing bit precision [67, 68]. GPTQ [69] leverages second-order information for error compensation, significantly enhancing model accuracy and training efficiency. The authors of [70] propose an edge-cutting quantization method, named AWQ, to quantize weights into low-bit integers, thereby reducing the hardware barrier and speeding up token generation. Unlike previous work on quantization for LLM inference, SwitchBack layers [71] investigates backpropagation through quantized weights over 1B parameters and QLoRA [62] quantizes weights into 4-bit NormalFloat and backpropagates gradients into low-rank structure with the pre-trained LLMs.
@@
Resource heterogeneity: The resource heterogeneity [50, 29, 72] poses a significant challenge to deploying FL frameworks [73, 74, 75] for LLMs. Clients with limited computing or memory resources require more time to complete local model updates, impeding the FL training process or even being excluded from model training. Although there are papers considering resource heterogeneity in general FL [76], there is no existing work that addresses resource heterogeneity in FL frameworks for LLM fine-tuning.
VII Conclusion
@@
In this paper, we have proposed and implemented FedPipe, an automated federated pipeline, to facilitate LLM FL fine-tuning in edge servers with heterogeneous computing resources. We have modeled the problem as a MILP to guide our design. To solve the problem, we have first selected important weights, and configured the corresponding parameters under computing budgets at edge servers for their LoRA adapters. We have also quantized LoRA adapters to further reduce memory space. Our extensive evaluations have demonstrated that FedPipe achieves significantly better performance than the state-of-the-art baselines in LLM FL.